OpenStudy (anonymous):

i have a question. for linear equations AFTER I SOLVE it in like point slope or like y intercept form and ext. how do i put my answer back in general form..

OpenStudy (phi):

can you give an example of the problem?

OpenStudy (anonymous):

idk

OpenStudy (phi):

You don't know your question?

OpenStudy (anonymous):

yes i do can you give me a second

OpenStudy (anonymous):

write an equation of the line through (5,5) and parallel to the line 4x+3y+-2

OpenStudy (phi):

typo in the equation? can you fix it?

OpenStudy (anonymous):

yeaj sorry it is and equal = not +

OpenStudy (anonymous):

so what i do is put it in y=mx+b so i can plug it in the calc. so after in that form what do i do

OpenStudy (phi):

4x+3y= -2 ? I would put this in the form y= mx+b so you can find the slope (m) can you do that ? the first step is write -4x+ on both sides of the = sign.

OpenStudy (anonymous):

no i got the ansewr in y=mx+b but the answer in the box gives it in general form

OpenStudy (anonymous):

and i dont have my answer with me so i can tell you

OpenStudy (phi):

what do you get after you put 4x+3y= -2 in y= mx+b form?

OpenStudy (anonymous):

y=-4x-2/3

OpenStudy (anonymous):

the whole equation over 3

OpenStudy (anonymous):

except y

OpenStudy (phi):

if you wrote it as y= (-4/3)x-2/3 it would be better. so the slope is -4/3 if you sub in (5,5) into y = mx +b with m= -4/3 (parallel to the line they gave you) 5= -4*5/3 +b to find b did you find b?

OpenStudy (anonymous):

yeah plug in 5 for y and x

OpenStudy (phi):

so the equation they want (in slope-intercept form) is y= (-4/3) x +35/3 is that what you got?

OpenStudy (anonymous):

yes but they want it in general form i put the answer in intercept form so i can plug it in the calc

OpenStudy (phi):

standard form is a x + by = c where a is positive (and not a fraction) add +(4/3)x to both sides then multiply both sides, all terms by 3

OpenStudy (anonymous):

ax+by=c or ax+by+c=0

OpenStudy (phi):

See http://en.wikipedia.org/wiki/Equation_of_a_line#Forms_for_2D_linear_equations for definitions of the different versions

OpenStudy (phi):

ax+by+c=0 is general form

OpenStudy (phi):

y= (-4/3) x +35/3 (4/3)x + y = 35/3 4x+3y= 35